OpenStudy (anonymous):

2x−3 over x−7 − x+1 over x−7 subtracting fraction, help :(

OpenStudy (campbell_st):

well you have a common denominator of x - 7 so the only danger... is getting the signs correct \[\frac{2x -3}{x -7} - \frac{(x + 1)}{x - 7} = \frac{2x -7 -(x +1)}{x - 7}\] just simplify the numerator ... carefully

OpenStudy (anonymous):

i got it, thank you ;3